A horse and a dog are running with the same speed. If the weight of the horse is 10 times that of the dog, what is the ratio of their kinetic energy?
1
Expert's answer
2013-01-29T10:37:06-0500
Let the weight of a dog be m. Then the horse weight is 10m. If they are running with the same speed V, the kinetic energy of a dog is

K(dog) = mV²/2

and the kinetic energy of a horse is

K(horse) = 10mV²/2.

Therefore, the ratio of their kinetic energy is

R = K(dog)/K(horse) = (mV²/2) / (10mV²/2) = 0.1.